OpenStudy (anonymous):

Can someone please help me? A child is creating a pyramid with building blocks. The top three levels include 3 blocks, 7 blocks, and 11 blocks. Part 1: How many blocks would be needed for a pyramid 25 levels tall? (5 points) Part 2: Use complete sentences to explain how a sum of an arithmetic series was applied. (4 points)

OpenStudy (ivettef365):

what is the sequence

OpenStudy (anonymous):

3,7,11, find the difference between each number

OpenStudy (anonymous):

it goes by +4

OpenStudy (anonymous):

yes, so your d in formula is equal to 4 formula Sum=a+(n-1)d

OpenStudy (anonymous):

where a=the first number in the sequence, in this case is 3

OpenStudy (anonymous):

and n is the nth term, and again in this case is 25

OpenStudy (anonymous):

so it would be 4=3+(25-1)(4)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Sum=3+(25-1)4

OpenStudy (anonymous):

99 blocks

OpenStudy (anonymous):

ooohh k thanks